OpenStudy (anonymous):

What is an equation of the tangent line at x = 5, assuming that y(5) = 4 and y'(5) = 2?

OpenStudy (freckles):

We have all the information we need to fill in the blanks. No real math is needed. Recall the equation of a tangent line is a line and therefore can be written as y-y1=m(x-x1) Where m is the slope at x=x1 and (x1,y1) is a point on the curve

OpenStudy (anonymous):

So it would be the points (5,4) (5,2)?

OpenStudy (freckles):

I only see on point on y that is (5,4)

OpenStudy (freckles):

The other equation you wrote is giving you the slope at x=5

OpenStudy (freckles):

That little ' denotes derivative (slope)

OpenStudy (anonymous):

Oh! so the slope is 2? so it would be y-4=(2(x-5)

OpenStudy (freckles):

looks good

OpenStudy (anonymous):

wow thank you so much for your help!!